I need to solve this algebra problem 6x+7=13x-9
Answer by Shana-D77(132)

6x+7=13x-9
Step 1: Find the =
Step 2: Find the + or -. What side is it on?
Since it's on both sides, pick your favorite term and add or subtract it. My favorite term is -9, so I will add 9 to both sides:
6x+7=13x-9
6x + 16 = 13x (added 9 to both sides)
Step 3: Find the + or - again. What side is it on? Since it's on the left, ask "Which term, 6x or 16, os a like term to the 13x on the right?" ANSWER: 6x
So, we subtract 6x from both sides:
6x + 16 = 13x
16 = 7x (subtracted 6x from both sides)
x = 16/7 (divided both sides by 7)